Winning Percentage. In a basketball league, the Falcons won 15 of their first 20 games. In order to win 60% of the total number of games, how many more games will they have to play, assuming they win only half of the remaining games?
I was able to come up with the answer of 50 games and 30 wins is 60% but to translate it is where I'm having a problem. Please help me.
Thank you, Nana

Answer by sudhanshu_kmr(1152)


Let x games have to play.
total no. of games = x + 20
no. of winning games = 15 + x/2

according to question,
15 + x/2 = 60 % of (x+20)
=> 15 + x/2 = (x+20) * 60/100
=> 15 + x/2 = 60x/100 + 12
=> 15 - 12 = 60x/100- x/2
=> 3 = x/10
=> x = 30
so, total 30 games have to play.......